OpenStudy (anonymous):

I am confused on this problem please help i tried it and i get something different each time: Justify each step by identifying the property used: 1/2(y+3)>1/3(4-y) 3(y+3)>2(4-y) 3y+9>8-2y 5y+9>8 5y>-1 y>-0.2

hero (hero):

I don't see anything wrong with it. Just don't forget to identify the properties.

OpenStudy (anonymous):

\[y>-\frac{1}{5} \]

hero (hero):

No need to state the obvious @robtobey \(y>-\frac{1}{5} ; -1/5 = -0.2\)